# revtex4-2Repair the float

### v2

@@ Acknowledgements
We acknowledge discussions with Alexander Golubov, Pauli Virtanen, Stefan Ilic, and Tero Heikkilä. T.K. and F.S.B. acknowledge financial support from Spanish MCIN/AEI/ 10.13039/501100011033 through project PID2020-114252GB-I00 (SPIRIT) and TED2021-130292B-C42, and the Basque Government through grant IT-1591-22. I.V.T. acknowledges support by Grupos Consolidados UPV/EHU del Gobierno Vasco (Grant IT1453-22) and by the grant PID2020-112811GB-I00 funded by MCIN/AEI/10.13039/501100011033.
@@ Appendix A Charge conjugation symmetry
In the theoretical description of superconductivity, it is convenient to introduce the bispinor of Fermion fields $\Psi=(\psi_{\uparrow},\psi_{\downarrow},\psi^{\dagger}_{\downarrow},-\psi^{\dagger}_{\uparrow})^{T}$ [25, 29, 30]. The bispinor $\Psi$ contains both $(\psi_{\uparrow},\psi_{\downarrow})$ and its time-reversed partner $(\psi_{\downarrow}^{\dagger},-\psi_{\uparrow}^{\dagger})$, where the time reversal operator is defined as $\sigma_{2}\mathcal{K}$. Here $\sigma_{2}$ is the second Pauli matrix in spin space, and $\mathcal{K}$ denotes complex conjugation. This is suitable for the description of superconductivity, which couples time-reversed pairs of electrons [53]. The theory is most conveniently described in terms of $\Psi$ and $\bar{\Psi}=\psi^{\dagger}\tau_{3}=(\psi_{\uparrow}^{\dagger},\psi_{\downarrow}^{\dagger},-\psi_{\downarrow},\psi_{\uparrow})$ instead of $\Psi$ and $\Psi^{\dagger}$, so that any term that is the same for electrons and holes comes along with a $\tau_{0}$, while any term that is opposite for electrons and holes comes with $\tau_{3}$.
@@
Since the Q-matrix is used for Hubbard-Stratonovich decoupling of $\Psi\bar{\Psi}$ [29, 31], the domain of integration for the Q-matrix should be restricted to those Q-matrices that satisfy this same symmetry. In contrast to the main text, here we distinguish the Q-matrices on the contour from those written as matrices in Keldysh space by using the calligraphic $\mathcal{Q}$ on the contour and a non-calligraphic $Q$ in Keldysh space. On the contour the charge conjugation symmetry can be written in the same notation as the symmetry on $\Psi\bar{\Psi}$:
$\mathcal{Q}_{t_{1},t_{2}}=C\mathcal{Q}_{t_{2},t_{1}}^{T}C^{-1},$ (13)
@@
In the main text the non-linear sigma model is, for convenience, presented in the rotated Keldysh space. First, the non-rotated Keldysh space is created by writing the different branches in matrix form, that is,
$Q^{\prime}_{t,t^{\prime}}=\begin{bmatrix}\mathcal{Q}_{t^{+},t^{\prime+}}&\mathcal{Q}_{t^{+},t^{\prime-}}\\ \mathcal{Q}_{t^{-},t^{\prime+}}&\mathcal{Q}_{t^{-},t^{\prime-}},\end{bmatrix}$ (14)
@@
$Q^{\prime}_{t,t^{\prime}}=CQ_{t,t^{\prime}}^{\prime T}C^{-1}.$ (15)
@@
$Q_{t,t^{\prime}}=L\rho_{3}Q_{t,t^{\prime}}^{\prime}L^{\dagger},$ (16)
@@
$Q_{t,t^{\prime}}=L\rho_{3}\tilde{Q}_{t,t^{\prime}}L^{-1}=L\rho_{3}C\tilde{Q}_{t,t^{\prime}}^{T}C^{-1}L^{-1}=L\rho_{3}CL^{-1}Q_{t,t^{\prime}}^{T}L\rho_{3}C^{-1}L^{-1}=L\rho_{3}L^{-1}CQ_{t,t^{\prime}}^{T}C^{-1}L\rho_{3}L^{-1},$ (17)
@@
$Q_{t,t^{\prime}}=\rho_{1}CQ_{t,t^{\prime}}^{T}C\rho_{1}=\rho_{1}\tau_{1}\sigma_{2}Q_{t,t^{\prime}}^{T}\sigma_{2}\tau_{1}\rho_{1}.$ (18)
This symmetry is referred to as the charge conjugation symmetry and introduced in Eq. 2 in the main text.
@@ Appendix B Chronological symmetry
  • The contour ordered Green’s function $i\mathcal{G}_{C}=\langle\mathcal{T}_{C}\Psi^{\dagger}(t_{1})\Psi(t_{2})\rangle$, where $\mathcal{T}_{C}$ denotes contour ordering on $C$ can be transformed into the anti-contour ordered function $i\mathcal{G}_{-C}=\langle\mathcal{T}_{-C}\Psi^{\dagger}(t_{1})\Psi(t_{2})\rangle$, where $\mathcal{T}_{-C}$ denotes anti-contour ordering on $C$ as can be obtained via Hermitian conjugation and adding a minus sign.
The latter condition is motivated by the generator of the time-evolution, the Hamiltonian, being Hermitian, in combination with the permutation rules for Grassmann variables [56, 57]. Continuity for $t_{1}\neq t_{2}$ is a consequence of the adiabatic assumption underlying the non-linear sigma model formalism, while the possibility of a discontinuity at $t_{1}=t_{2}$ arises because of the difference in order of the operators in the expressions of $\mathcal{G}_{C}$ for $t_{1}<t_{2}$ and $t_{1}>t_{2}$.
@@
$i\mathcal{G}_{-C}(t_{1},t_{2})=i\mathcal{G}^{<}(t_{1},t_{2})\theta_{C}(t_{1},t_{2})+i\mathcal{G}^{>}(t_{1},t_{2})\theta_{C}(t_{2},t_{1}),$ (19)
@@
$i\mathcal{G}_{-C}=-(i\mathcal{G}_{C})^{\dagger},$ (20)
$i\mathcal{G}^{>}=-(i\mathcal{G}^{<})^{\dagger}.$ (21)
@@
$i\mathcal{G}^{\tilde{\mathcal{T}}}=-(i\mathcal{G}^{\mathcal{T}})^{\dagger}.$ (22)
Next, we may write the contour ordered Green’s function $\mathcal{G}_{C}$ in the non-rotated Keldysh space as, where, like in Sec. A we distinguish in notation between contour representation and Keldysh space representation,
$iG^{\prime}(t,t^{\prime})=\begin{bmatrix}i\mathcal{G}_{C}(t^{+},t^{+})&i\mathcal{G}_{C}(t^{+},t^{-})\\ i\mathcal{G}_{C}(t^{-},t^{+})&i\mathcal{G}_{C}(t^{-},t^{-})\end{bmatrix}=\begin{bmatrix}i\mathcal{G}^{\mathcal{T}}(t,t^{\prime})&i\mathcal{G}^{>}(t,t^{\prime})\\ i\mathcal{G}^{<}(t,t^{\prime})&i\mathcal{G}^{\tilde{\mathcal{T}}}(t,t^{\prime})\end{bmatrix}.$ (23)
@@
$iG^{\prime}=-\rho_{1}(iG)^{\prime\dagger}\rho_{1}.$ (24)
@@
$=-\rho_{2}(L\rho_{3}(iG)^{\prime}L^{-1})^{\dagger}\rho_{2}=-\rho_{2}(iG)^{\dagger}\rho_{2}.$ (25)
@@ Appendix C Terms combining Pauli matrices and derivatives
In this section, we derive the different terms leading to Eqs. (4-6) in the main text. For this, we assume a weak SOC and determine the allowed terms in the action to second and third order in $\sigma$-matrices and derivatives.
@@ C.1 Second order terms
Terms containing two Pauli matrices or two derivatives, provide the general relaxation, $\Gamma_{ij}$, and diffusion tensors, $D_{ij}$, in Eq. (4). Here we focus on the derivation of other terms that contain one Pauli matrix and one derivative. There are two of such terms, $\sigma_{k}\partial_{j}Q$ and $\sigma_{k}Q\partial_{j}Q$. Applying transposition to the first we find
$\text{Tr}-\partial_{j}Q^{T}\sigma_{y}\sigma_{k}\sigma_{y}=-\text{Tr}\rho_{1}\tau_{1}\sigma_{y}\partial_{j}Q^{T}\sigma_{y}\rho_{1}\tau_{1}\sigma_{k}=-\text{Tr}\partial_{j}Q\sigma_{k}=-\text{Tr}\sigma_{k}\partial_{j}Q.$ (28)
To obtain a charge conjugation symmetric term, this term should be equivalent to the one we started from, hence this imposes the restriction $\alpha_{kj}=-\alpha_{kj}$ on the corresponding tensor, i.e. $\alpha_{kj}=0$. We conclude that this term is not allowed in the action.
@@
$\partial_{j}Q^{T}Q^{T}\sigma_{y}\sigma_{k}\sigma_{y}=-\text{Tr}\rho_{1}\tau_{1}\sigma_{y}\partial_{j}Q^{T}\sigma_{y}\rho_{1}\tau_{1}\sigma_{y}\rho_{1}\tau_{1}Q^{T}\tau_{1}\rho_{1}\sigma_{y}\sigma_{k}=-\text{Tr}\partial_{j}QQ\sigma_{k}=\text{Tr}\sigma_{k}Q\partial_{j}Q.$ (29)
@@
$\text{Tr}\partial_{j}Q^{\dagger}Q^{\dagger}\sigma_{k}=\text{Tr}\tau_{3}\rho_{2}\partial_{j}Q^{\dagger}\rho_{2}\tau_{3}\tau_{3}\rho_{2}Q^{\dagger}\rho_{2}\tau_{3}\sigma_{k}=\text{Tr}\partial_{j}\tilde{Q}\tilde{Q}\sigma_{k}=-\text{Tr}\sigma_{k}\tilde{Q}\partial_{j}\tilde{Q}$ (30)
where $\tilde{Q}=-\rho_{2}\tau_{3}Q^{\dagger}\tau_{3}\rho_{2}$. Since $iS(Q)^{*}=iS(\tilde{Q})$ it now follows that $\tilde{\alpha}_{kj}=-\tilde{\alpha}_{kj}^{*}$, that is, the tensor has only imaginary indices. Thus, $\tilde{\alpha}_{kj}=i\alpha_{kj}$, where $\alpha_{kj}$ is a real second rank pseudotensor. The corresponding term in the bulk equation, which can be derived via the requirement that at the saddle point the action is invariant under variations $g+[\delta g,g]$, is
$i\alpha_{kj}[[\sigma_{k},\partial_{j}g],g].$ (31)
In the Keldysh equation of the normal state this becomes $2i\alpha_{kj}[\sigma_{k},\partial_{j}F]$. Because the action contains terms with derivatives, integration by parts is necessary to derive the saddle point condition. This additionally yields terms that are evaluated only at the boundary. These terms have the form $n_{j}J_{j}$, where $n_{j}$ is the normal vector of the boundary and $J_{j}$ is the contribution to the current of the term under consideration. Equivalently, the expression for the current can be found by variation of the action with the vector potential. The contribution of the term $i\alpha_{kj}\sigma_{k}Q\partial_{j}Q$ to the current is
$J_{j}=i\alpha_{kj}[\sigma_{k}g,g].$ (32)
In the normal state the Keldysh component of this expression vanishes.
@@ C.2 Third order terms with two Pauli matrices and one derivative
There exist two terms of this order: $\sigma_{i}Q\sigma_{j}\partial_{k}Q$ and $\sigma_{i}Q\sigma_{j}Q\partial_{k}Q$. Transposition of the first term gives
$\text{Tr}\partial_{k}Q^{T}\sigma_{y}\sigma_{j}\sigma_{y}Q^{T}\sigma_{y}\sigma_{i}\sigma_{y}=\text{Tr}\rho_{1}\tau_{1}\sigma_{y}\partial_{k}Q^{T}\sigma_{y}\rho_{1}\tau_{1}\sigma_{j}\tau_{1}\rho_{1}\sigma_{y}Q^{T}\sigma_{y}\tau_{1}\rho_{1}\sigma_{i}=\text{Tr}\partial_{k}Q\sigma_{j}Q\sigma_{i}=\text{Tr}\sigma_{j}Q\sigma_{i}\partial_{k}Q,$ (33)
@@
$\text{Tr}\partial_{k}Q^{\dagger}\sigma_{j}Q^{\dagger}\sigma_{i}=\text{Tr}\rho_{2}\tau_{3}\partial_{k}Q^{\dagger}\rho_{2}\tau_{3}\sigma_{j}\rho_{2}\tau_{3}Q^{\dagger}\rho_{2}\tau_{3}\sigma_{i}=\text{Tr}\partial_{k}\tilde{Q}\sigma_{j}\tilde{Q}\sigma_{i}=\text{Tr}\sigma_{j}\tilde{Q}\sigma_{i}\partial_{k}Q,$ (34)
that is $\beta_{ijk}=\beta_{jik}^{*}=\beta_{ijk}$, where in the last equality we used the charge conjugation symmetry relation derived above. Thus, this tensor is real. The contribution to the bulk equation is
$\beta_{ijk}[\sigma_{i}\partial_{k}g\sigma_{j}-\sigma_{j}\partial_{k}g\sigma_{i},g]=0$ (35)
by symmetry in the $i,j$ indices. Indeed, we may write this term as a total derivative term $\partial_{k}(\sigma_{i}Q\sigma_{j}Q)$, and hence this term can only enter via the boundaries.
@@
$J_{k}=\beta_{ijk}[\sigma_{j}g\sigma_{i},g].$ (36)
In the normal state the Keldysh component is
$J_{k}^{K}=2\beta_{ijk}(\{F,\sigma_{j}\sigma_{i}\}-2\sigma_{j}F\sigma_{i}).$ (37)
By symmetry of $\beta_{ijk}$ upon exchange of $i,j$ we may write this as
$J_{k}^{K}=4\beta_{ijk}(F-\sigma_{j}F\sigma_{i})$ (38)
@@
$=\text{Tr}\partial_{k}QQ\sigma_{j}Q\sigma_{i}=-\text{Tr}\sigma_{j}Q\sigma_{i}Q\partial_{k}Q,$ (39)
that is, the third rank pseudotensor corresponding to this term is antisymmetric in the first two indices and we may write it as $\varepsilon_{ijl}\tilde{\gamma}_{lk}$, where $\varepsilon_{ijl}$ is the fully antisymmetric third rank pseudotensor and $\tilde{\gamma}_{lk}$ is a second rank pseudotensor.

### v3

@@ Acknowledgements
We acknowledge discussions with Alexander Golubov, Pauli Virtanen, Stefan Ilic and Tero Heikkilä. We acknowledge financial support from Spanish MCIN/AEI/ 10.13039/501100011033 through projects PID2023-148225NB-C31, PID2023-148225NB-C32(SUNRISE), and TED2021-130292B-C42, the Basque Government through grants IT-1591-22 and IT1453-22, and the European Union’s Horizon Europe research and innovation program under grant agreement No 101130224 (JOSEPHINE).
@@ Appendix A Charge conjugation symmetry
In the theoretical description of superconductivity, it is convenient to introduce the bispinor of Fermion fields $\Psi=(\psi_{\uparrow},\psi_{\downarrow},\psi^{\dagger}_{\downarrow},-\psi^{\dagger}_{\uparrow})^{T}$ [26, 31, 32]. The bispinor $\Psi$ contains both $(\psi_{\uparrow},\psi_{\downarrow})$ and its time-reversed partner $(\psi_{\downarrow}^{\dagger},-\psi_{\uparrow}^{\dagger})$, where the time reversal operator is defined as $\sigma_{2}\mathcal{K}$. Complex conjugation is denoted by $\mathcal{K}$ and $\sigma_{2}$ is the second Pauli matrix in spin space. This is suitable for the description of superconductivity, which couples time-reversed pairs of electrons [41]. The theory is most conveniently described in terms of $\Psi$ and $\bar{\Psi}=\psi^{\dagger}\tau_{3}=(\psi_{\uparrow}^{\dagger},\psi_{\downarrow}^{\dagger},-\psi_{\downarrow},\psi_{\uparrow})$ instead of $\Psi$ and $\Psi^{\dagger}$, so that any term that is the same for electrons and holes comes along with a $\tau_{0}$, while any term that is opposite for electrons and holes comes with $\tau_{3}$.
@@
Since the Q-matrix is used for Hubbard-Stratonovich decoupling of $\Psi\bar{\Psi}$ [31, 33], the domain of integration for the Q-matrix should be restricted to those Q-matrices that satisfy this same symmetry. In contrast to the main text, here we distinguish the Q-matrices on the contour from those written as matrices in Keldysh space by using the calligraphic $\mathcal{Q}$ on the contour and a noncalligraphic $Q$ in Keldysh space. On the contour the charge conjugation symmetry can be written in the same notation as the symmetry on $\Psi\bar{\Psi}$:
$\mathcal{Q}_{t_{1},t_{2}}=C\mathcal{Q}_{t_{2},t_{1}}^{T}C^{-1}\;,$ (S13)
@@
In the main text the nonlinear sigma model is, for convenience, presented in the rotated Keldysh space. First, the nonrotated Keldysh space is created by writing the different branches in matrix form, that is,
$Q^{\prime}_{t,t^{\prime}}=\begin{bmatrix}\mathcal{Q}_{t^{+},t^{\prime+}}&\mathcal{Q}_{t^{+},t^{\prime-}}\\ \mathcal{Q}_{t^{-},t^{\prime+}}&\mathcal{Q}_{t^{-},t^{\prime-}}\;,\end{bmatrix}$ (S14)
@@
$Q^{\prime}_{t,t^{\prime}}=CQ_{t,t^{\prime}}^{\prime T}C^{-1}\;.$ (S15)
@@
$Q_{t,t^{\prime}}=L\rho_{3}Q_{t,t^{\prime}}^{\prime}L^{\dagger}\;,$ (S16)
@@
$Q_{t,t^{\prime}}=L\rho_{3}\tilde{Q}_{t,t^{\prime}}L^{-1}=L\rho_{3}C\tilde{Q}_{t,t^{\prime}}^{T}C^{-1}L^{-1}=L\rho_{3}CL^{-1}Q_{t,t^{\prime}}^{T}L\rho_{3}C^{-1}L^{-1}=L\rho_{3}L^{-1}CQ_{t,t^{\prime}}^{T}C^{-1}L\rho_{3}L^{-1}\;,$ (S17)
@@
$Q_{t,t^{\prime}}=\rho_{1}CQ_{t,t^{\prime}}^{T}C\rho_{1}=\rho_{1}\tau_{1}\sigma_{2}Q_{t,t^{\prime}}^{T}\sigma_{2}\tau_{1}\rho_{1}\;.$ (S18)
This symmetry is referred to as the charge conjugation symmetry and introduced in Eq. (2) in the main text.
@@ Appendix B Chronological symmetry
  • The contour ordered Green’s function $i\mathcal{G}_{C}=\langle\mathcal{T}_{C}\Psi^{\dagger}(t_{1})\Psi(t_{2})\rangle$, where $\mathcal{T}_{C}$ denotes contour ordering on $C$ can be transformed into the anti-contour ordered function $i\mathcal{G}_{-C}=\langle\mathcal{T}_{-C}\Psi^{\dagger}(t_{1})\Psi(t_{2})\rangle$, where $\mathcal{T}_{-C}$ denotes anti-contour ordering on $C$ via Hermitian conjugation and adding a minus sign.
The latter condition is motivated by the generator of the time-evolution, the Hamiltonian, being Hermitian, in combination with the permutation rules for Grassmann variables [30, 36]. Continuity for $t_{1}\neq t_{2}$ is a consequence of the adiabatic assumption underlying the nonlinear sigma model formalism, while the possibility of a discontinuity at $t_{1}=t_{2}$ arises because of the difference in order of the operators in the expressions of $\mathcal{G}_{C}$ for $t_{1}<t_{2}$ and $t_{1}>t_{2}$.
@@
$i\mathcal{G}_{-C}(t_{1},t_{2})=i\mathcal{G}^{<}(t_{1},t_{2})\theta_{C}(t_{1},t_{2})+i\mathcal{G}^{>}(t_{1},t_{2})\theta_{C}(t_{2},t_{1})\;,$ (S19)
@@
$i\mathcal{G}_{-C}=-(i\mathcal{G}_{C})^{\dagger}\;,$ (S20)
$i\mathcal{G}^{>}=-(i\mathcal{G}^{<})^{\dagger}\;.$ (S21)
@@
$i\mathcal{G}^{\tilde{\mathcal{T}}}=-(i\mathcal{G}^{\mathcal{T}})^{\dagger}\;.$ (S22)
Next, we may write the contour ordered Green’s function $\mathcal{G}_{C}$ in the nonrotated Keldysh space as, where, like in Sec. A we distinguish in notation between contour representation and Keldysh space representation,
$iG^{\prime}(t,t^{\prime})=\begin{bmatrix}i\mathcal{G}_{C}(t^{+},t^{+})&i\mathcal{G}_{C}(t^{+},t^{-})\\ i\mathcal{G}_{C}(t^{-},t^{+})&i\mathcal{G}_{C}(t^{-},t^{-})\end{bmatrix}=\begin{bmatrix}i\mathcal{G}^{\mathcal{T}}(t,t^{\prime})&i\mathcal{G}^{>}(t,t^{\prime})\\ i\mathcal{G}^{<}(t,t^{\prime})&i\mathcal{G}^{\tilde{\mathcal{T}}}(t,t^{\prime})\end{bmatrix}\;.$ (S23)
@@
$iG^{\prime}=-\rho_{1}(iG)^{\prime\dagger}\rho_{1}\;.$ (S24)
@@
$=-\rho_{2}(L\rho_{3}(iG)^{\prime}L^{-1})^{\dagger}\rho_{2}=-\rho_{2}(iG)^{\dagger}\rho_{2}\;.$ (S25)
@@ Appendix C Terms combining Pauli matrices and derivatives
In this section, we derive the different terms leading to Eqs. (4-6) in the main text. For this, we determine the allowed terms in the action to second and third order in $\sigma$-matrices and derivatives.
@@ C.1 Second order terms
Terms containing two Pauli matrices or two derivatives, provide the general relaxation, $\Gamma_{ij}$, and diffusion tensors, $D_{ij}$, in Eq. (4) in the main text. Here we focus on the derivation of other terms that contain one Pauli matrix and one derivative. There are two of such terms,
$\bar{\alpha}_{kj}\sigma_{k}\partial_{j}Q\;,$ (S28)
and
$\frac{i}{2}\alpha_{kj}\sigma_{k}Q\partial_{j}Q\;.$ (S29)
Applying transposition to the first we find
$\text{Tr}-\partial_{j}Q^{T}\sigma_{y}\sigma_{k}\sigma_{y}=-\text{Tr}\rho_{1}\tau_{1}\sigma_{y}\partial_{j}Q^{T}\sigma_{y}\rho_{1}\tau_{1}\sigma_{k}=-\text{Tr}\partial_{j}Q\sigma_{k}=-\text{Tr}\sigma_{k}\partial_{j}Q\;.$ (S30)
To obtain a charge conjugation symmetric term, this term should be equivalent to the one we started from, hence this imposes the restriction $\bar{\alpha}_{kj}=-\bar{\alpha}_{kj}$ on the corresponding tensor, i.e. $\bar{\alpha}_{kj}=0$. We conclude that this term is not allowed in the action.
@@
$\text{Tr}\partial_{j}Q^{T}Q^{T}\sigma_{y}\sigma_{k}\sigma_{y}=-\text{Tr}\rho_{1}\tau_{1}\sigma_{y}\partial_{j}Q^{T}\sigma_{y}\rho_{1}\tau_{1}\sigma_{y}\rho_{1}\tau_{1}Q^{T}\tau_{1}\rho_{1}\sigma_{y}\sigma_{k}=-\text{Tr}\partial_{j}QQ\sigma_{k}=\text{Tr}\sigma_{k}Q\partial_{j}Q\;.$ (S31)
@@
$\text{Tr}\partial_{j}Q^{\dagger}Q^{\dagger}\sigma_{k}=\text{Tr}\tau_{3}\rho_{2}\partial_{j}Q^{\dagger}\rho_{2}\tau_{3}\tau_{3}\rho_{2}Q^{\dagger}\rho_{2}\tau_{3}\sigma_{k}=\text{Tr}\partial_{j}\tilde{Q}\tilde{Q}\sigma_{k}=-\text{Tr}\sigma_{k}\tilde{Q}\partial_{j}\tilde{Q}\;,$ (S32)
where $\tilde{Q}=-\rho_{2}\tau_{3}Q^{\dagger}\tau_{3}\rho_{2}$. Since $iS(Q)^{*}=iS(\tilde{Q})$ it now follows that $\alpha_{kj}=\alpha_{kj}^{*}$, that is, $\alpha_{kj}$ is a real second rank pseudotensor. The corresponding term in the bulk equation, which can be derived via the requirement that at the saddle point the action is invariant under variations $g+[\delta g,g]$:
$i\text{Tr}\Big{(}\alpha_{kj}\sigma_{k}[\delta g,g]\partial_{j}g+\alpha_{kj}\sigma_{k}g\partial_{j}[\delta g,g]\Big{)}\;.$ (S33)
If we use the cyclic property of the trace to write $\delta g$ and the end, we obtain
$-\text{Tr}\Bigg{(}i\alpha_{kj}\Big{(}\sigma_{k}\partial_{j}g\sigma_{k}g-g\partial_{j}g\sigma_{k}+\sigma_{k}g\partial_{j}g-\partial_{j}g\sigma_{k}g\Big{)}\delta g+i\alpha_{kj}\Big{(}\sigma_{k}-g\sigma_{k}g\Big{)}\partial_{j}\delta g\Bigg{)}\;.$ (S34)
After doing integration by parts, we obtain the bulk contribution
$-\text{Tr}\Big{(}i\alpha_{kj}[\sigma_{k},g\partial_{j}g]+\partial_{j}(i\alpha_{kj}g[\sigma_{k},g])\Big{)}\delta g\;.$ (S35)
It is convenient to separate this contribution into the current and the torque. The current may be obtained via variation of the action with the vector potential or equivalently by evaluation of the boundary terms that arise upon the required integration by parts in the derivation of the saddle point. The contribution of the term $i\alpha_{kj}\sigma_{k}Q\partial_{j}Q$ to the current is
$J_{j}=i\alpha_{kj}g[\sigma_{k},g]\;.$ (S36)
Consequently, the torque reads
$\mathcal{T}=-i\alpha_{kj}[\sigma_{k},g\partial_{j}g]\;.$ (S37)
@@ C.2 Third order terms with two Pauli matrices and one derivative
There exist two terms of this order:
$\beta_{ijk}\sigma_{i}Q\sigma_{j}\partial_{k}Q$ (S38)
and
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-}\frac{i}{8}\gamma_{ijk}\sigma_{i}Q\sigma_{j}Q\partial_{k}Q$ (S39)
. Transposition of the first term gives
$\text{Tr}\partial_{k}Q^{T}\sigma_{y}\sigma_{j}\sigma_{y}Q^{T}\sigma_{y}\sigma_{i}\sigma_{y}=\text{Tr}\rho_{1}\tau_{1}\sigma_{y}\partial_{k}Q^{T}\sigma_{y}\rho_{1}\tau_{1}\sigma_{j}\tau_{1}\rho_{1}\sigma_{y}Q^{T}\sigma_{y}\tau_{1}\rho_{1}\sigma_{i}=\text{Tr}\partial_{k}Q\sigma_{j}Q\sigma_{i}=\text{Tr}\sigma_{j}Q\sigma_{i}\partial_{k}Q\;,$ (S40)
@@
$\text{Tr}\partial_{k}Q^{\dagger}\sigma_{j}Q^{\dagger}\sigma_{i}=\text{Tr}\rho_{2}\tau_{3}\partial_{k}Q^{\dagger}\rho_{2}\tau_{3}\sigma_{j}\rho_{2}\tau_{3}Q^{\dagger}\rho_{2}\tau_{3}\sigma_{i}=\text{Tr}\partial_{k}\tilde{Q}\sigma_{j}\tilde{Q}\sigma_{i}=\text{Tr}\sigma_{j}\tilde{Q}\sigma_{i}\partial_{k}Q\;,$ (S41)
that is $\beta_{ijk}=\beta_{jik}^{*}=\beta_{ijk}$, where in the last equality we used the charge conjugation symmetry relation derived above. Thus, this tensor is real. The contribution to the bulk equation can be found using variation:
$\beta_{ijk}\sigma_{i}[\delta g,g]\sigma_{j}\partial_{k}g+\sigma_{i}g\sigma_{j}\partial_{k}[\delta g,g]$ (S42)
$=\beta_{ijk}\Big{(}g\sigma_{j}\partial_{k}g\sigma_{i}-\sigma_{j}\partial_{k}g\sigma_{i}g+\partial_{k}g\sigma_{i}g\sigma_{j}-\sigma_{i}g\sigma_{j}\partial_{k}g\Big{)}\delta g+\Big{(}g\sigma_{i}g\sigma_{j}-\sigma_{i}g\sigma_{j}g\Big{)}\partial_{k}g$ (S43)
$=\Bigg{(}-\beta_{ijk}\partial_{k}[\sigma_{i}g\sigma_{j},g]+\partial_{k}(\beta_{ijk}[\sigma_{i}g\sigma_{j},g])\Bigg{)}\delta g\;.$ (S44)
Specifically, if $\beta_{ijk}$ is constant, this reduces to
$0\;.$ (S45)
by symmetry in the $i,j$ indices. Indeed, in that case we may write this term as a total derivative term $\partial_{k}(\sigma_{i}Q\sigma_{j}Q)$, and hence this term can only enter via the boundaries.
@@
$J_{k}=-\beta_{ijk}[\sigma_{j}g\sigma_{i},g]\;.$ (S46)
Consequently, the torque is given by
$\mathcal{T}=-\beta_{ijk}\partial_{k}[\sigma_{j}g\sigma_{i},g]\;.$ (S47)
@@
$=\text{Tr}\partial_{k}QQ\sigma_{j}Q\sigma_{i}=-\text{Tr}\sigma_{j}Q\sigma_{i}Q\partial_{k}Q\;,$ (S48)
that is, the third rank pseudotensor corresponding to this term is antisymmetric in the first two indices and we may write it as $\frac{1}{8}\varepsilon_{ijl}\gamma_{lk}$, where $\varepsilon_{ijl}$ is the fully antisymmetric third rank pseudotensor and $\tilde{\gamma}_{lk}$ is a second rank pseudotensor. The $\frac{1}{8}$ is chosen for convenience.